The mean of the times it takes a commuter to get to work in Baltimore is 29.7 minutes. If the standard deviation is 6 minutes, within what liits would you expect approximately 68% of the times to fall? Assume the distribution is approximately bell-shaped.
Found 2 solutions by stanbon, drk:
Answer by stanbon(75887)

The mean of the times it takes a commuter to get to work in Baltimore is 29.7 minutes. If the standard deviation is 6 minutes, within what liits would you expect approximately 68% of the times to fall? Assume the distribution is approximately bell-shaped.
---------------
68% of the times are within one standard deviation of the mean time.
-----
mean-std = 29.7-6 = 23.7 minutes
---
mean+std = 29.7+6 = 35.7 minutes
====================================
Cheers,
Stan H.

Answer by drk(1908)

68% is one standard deviation above and one standard deviation below the mean.
So, we take 29.7 + and - 6 and we get
23.7 to 35.7 minutes in 68% of our data.